We will use the equation for the magnetic force be
Answer #1

First of all a figure must be attached with the question. The numerical values should be given clearly.

Here the force is given by,

\vec{F} = \vec{F}_1 + \vec{F}_2 = \frac{\mu_0 I_1 I_2 l}{2\pi} \left( \frac{1}{c+a} - \frac{1}{c}\right ) \hspace{2mm} \hat{i} = \frac{\mu_0 I_1 I_2 l}{2\pi} \left( - \frac{a}{c(c+a)}\right ) \hspace{2mm} \hat{i}

From the question, we can get the following information,

I_1 I_2 = 87 \hspace{2mm} A^2, \hspace{8mm} l = 0.595 \hspace{2mm } m, \hspace{8mm} a = 0.15 \hspace{2mm} m

c(c+a) = 0.025 \hspace{2mm} m^2

Using these values we get,

\vec{F} = \frac{4\pi \times 10^{-7} \times 87 \times 0.595}{2\pi} \times \left( - \frac{0.15}{0.025}\right ) \hspace{2mm} \hat{i} = -6.2 \times 10^{-5} \hspace{2mm} \hat{i} \hspace{2mm} \text{N}

Hence the magnitude of the total force is,

|\vec{F}| = 6.2 \times 10^{-5} \hspace{2mm} N = 62 \hspace{2mm} \mu N

And the force is directed towards the -x axis.